OpenStudy (anonymous):

f(x)=x^2-1 and g(x)=2x-3 find (g o f)(5) Help PLEASE :)

OpenStudy (anonymous):

Find f(5), plug the answer into g(x)

OpenStudy (anonymous):

what is f(5) if f(x)=x^2-1?

OpenStudy (anonymous):

(g o f)(x)=g(f(x))

OpenStudy (ash2326):

\[gof(x)=g(f(x))\] we have g(x)=2x-3 f(x)=x^2-1 so \[gof(x)=2(x^2-1)-3\] we get \[gof(x)=2x^2-5\] put x=5 \[gof(5)=2*5*5-5\] we get \[gof(5)=45\]

OpenStudy (anonymous):

45 or -45?

OpenStudy (ash2326):

45